Z ring of integers, R ring, f: ring homomorphism from Z to R which is injective. Also if n is not in 13Z, then f(n) is unit in R. Can R be integral domain. One example of such a map.
1
Expert's answer
2020-03-03T07:34:32-0500

The ring "R" can be an integral domain. Let "R" be "\mathbb{Q}" and "f" be defined by "f(n) = n". Obviously, "f" is injective and a ring homomorphism.


Assume that "n\not\in 13 \mathbb{Z}". Since "0\in 13 \mathbb{Z}", "n\neq 0". Then "f(n)\neq 0", so "f(n)" is a unit because "\mathbb{Q}" is a field. Every field is an integral domain, hence "R" is an integral domain.